OpenStudy (anonymous):

I need help with multiplying, dividing, adding, and subtracting radical expressions in algebra.

OpenStudy (anonymous):

Such as?

OpenStudy (anonymous):

just in general how to do it

OpenStudy (anonymous):

Post a specific question so it's easier to guide you through it.

OpenStudy (anonymous):

We are simplifying them in class

OpenStudy (anonymous):

sqrt(3)sqrt(10)

OpenStudy (anonymous):

\[\sqrt{3}\sqrt{10}\] Like the above?

OpenStudy (anonymous):



OpenStudy (anonymous):

ya but with a multiplication sign

OpenStudy (anonymous):

In this case, all you can do is multiply them together. When multiplying radicals, as long as they have the same index, then you can multiply the numbers inside the radicals. \[\sqrt{30}\]

OpenStudy (anonymous):

http://www.wtamu.edu/academic/anns/mps/math/mathlab/int_algebra/int_alg_tut39_simrad.htm#prod is a pretty good page for reviewing radicals.